OpenStudy (anonymous):

Some investments in the stock market have earned 10% annually. At this rate, earnings can be found using the formula A=P(1.10)^r, where A is the total value of the investment, P is the initial value of the investment, and n is the number of years the money is invested. If $1,500 is invested in the stock market at this annual rate of return, what is expected of the total value in 18 years?

OpenStudy (kasim17):

27000

OpenStudy (anonymous):

My choices are 29,7000.00, 28,050.00, 8,339.88 and 7, 581.71

OpenStudy (kasim17):

oh ok thought my answer was too big so i started over u should put those in beginning

OpenStudy (anonymous):

Sorry, I'm in a bit of a rush.

OpenStudy (kasim17):

finals?

OpenStudy (kasim17):

28,050.00 im not entirely sure though

OpenStudy (anonymous):

No, just a normal assignment. I'm just very behind.

OpenStudy (anonymous):

Thank you.

OpenStudy (kasim17):

np

OpenStudy (kasim17):

doesn't hurt to double check though with someone else

OpenStudy (wolf1728):

Total = Principal*(1+r)^n Total = 1,500.00 * (1+10)^18 Total = 1,500.00 * (5.5599173135) Total = 8,339.88 For a double-check, enter the numbers here: http://www.1728.org/compint.htm

OpenStudy (wolf1728):

Second line should read 1,500.00 * (1.10)^18

OpenStudy (anonymous):

Alright, thank you.

OpenStudy (wolf1728):

u r welcome

OpenStudy (anonymous):

I fanned one of you and best response the other, so it's equal. Thank you both.

OpenStudy (kasim17):

np